OpenStudy (anonymous):

Jared has a wooden board that is 11 ft long. He wants to cut it into two pieces so that the longer piece is 4 ft less than 4 times the length of the shorter piece. How long will each piece of board be? Shorter piece:____ft Longer piece:____ft

Directrix (directrix):

Let x and y represent the two pieces of board. So, x + y + ?

OpenStudy (anonymous):

x+y=11

Directrix (directrix):

Now, for the second equation.

Directrix (directrix):

longer piece is 4 ft less than 4 times the length of the shorter piece

Directrix (directrix):

It does not matter what variable is coded and longer or shorter. One of them is 4 times the other minus 4

Directrix (directrix):

If x is one of them, and y the other, then the equation is: x =

OpenStudy (anonymous):

x=4

Directrix (directrix):

Not that.

OpenStudy (anonymous):

what do you mean

Directrix (directrix):

longer piece is 4 ft less than 4 times the length of the shorter piece x = 4* ? - 4

OpenStudy (anonymous):

x=4y-4

OpenStudy (anonymous):

corret?

OpenStudy (anonymous):

brb have to let dog in

Directrix (directrix):

x=4y-4 x+y=11 We are taking these to the wolf at http://www.wolframalpha.com/ Enter as this: x=4y-4, x+y=11

Directrix (directrix):

What did you get?

OpenStudy (anonymous):

x=8 y=3

Directrix (directrix):

That is what I got. Which option? Close and post.

OpenStudy (anonymous):

shorter piece=8 longer piece=3

Directrix (directrix):

Shorter piece:____ft Longer piece:____ft Be sure to put the right number in the right slot

Directrix (directrix):

Look at this: shorter piece=8 longer piece=3

Directrix (directrix):

Is 8 shorter than 3? Keep your focus.

OpenStudy (anonymous):

sorry shorter piece=3 longer piece=8

OpenStudy (anonymous):

thanks